Jen Butler has been pricing​ Speed-Pass train fares for a group trip to New York. Three adults and four children must pay $87. Two adults and three children must pay $62. Find the price of the​ adult's ticket and the price of a​ child's ticket.

Answer by ewatrrr(24785)


Hi
2A + 3C = $62
3A + 4C = $87
6A + 9C = $186 |1ST EQ X 3
6A + 8C = $174 |2ND EQ X 2
6A + 9C = $186
-6A - 8C = - $174
Child = $12 per ticket
2A = $62 - $36 = $26 1ST EQ
2A = $ 26
Adult = $13 per ticket
CHECKING our answer*** 2ND EQ
$39 + $48 = $87 CHECKS!
Wish You the Best in your Studies.